2tan30°/1-tan^2 30° =
(A) cos 60° (B) sin 60° (C) tan 60° (D) sin 30°

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option C is correct.

Substitute the of tan 30° in the given equation
tan 30° = 1/√3

2tan30°/1-tan^2 30° =2(1/√3)/1-(1/√3)^2

= (2/√3)/(1-1/3)
= (2/√3)/(2/3)
= √3
= tan 60°
The value of the given equation is equivalent to tan 60°.

Thanks